OpenStudy (anonymous):

Find tan B and write as a simplified fraction. Use the triangle below to answer this question.

OpenStudy (anonymous):

Do you remember the little saying SOH CAH TOA?

OpenStudy (anonymous):

tan x = opposite side over adjacent side.

OpenStudy (anonymous):

15/36?

OpenStudy (raden):

36/15 = 12/5 is the answer

OpenStudy (anonymous):

Find sin A and write as a simplified fraction. Use the triangle below to answer this question.

OpenStudy (anonymous):

this is another question

OpenStudy (anonymous):

thank you for the first one

OpenStudy (raden):

find the value of x, first use phytagorean theorema : x^2 = 15^2+36^2 x^2 = 225 + 1296 x^2 = 1521 x = sqrt(1461) x = 39 now, we can calculate the value of sinA sinA = 15/x = 15/39 = 5/13

OpenStudy (raden):

opss... typo error above (problem my keyboard) x= sqrt(1521)